Problem with solution/question:
Solve 2L + 2W = P for L.
2L = P - 2w Subtract 2W from each side.
L = P - 2W
------ Divide by 2.
2


Note with solution-- L = P - 2W
------
2

Can be rewritten as-- L = P/2 - W.
Question: Did they just add 2 to the 2W? None-homework.

Answer by josgarithmetic(39613)

Your method is correct.
Your notation for the site is poor.

Done in pure text,
2L=P-2W
(1/2)2L=(1/2)(P-2W)
L=(P-2W)2

But you can also do the distrubution for the factor (1/2) and say L=P/2-W.